OpenStudy (anonymous):

Use an addition or subtraction formula to find the exact value of tan(285)

OpenStudy (anonymous):

Cant you type this in your calculator?

OpenStudy (anonymous):

Tan(285) according to my Calc is -1.221

OpenStudy (dumbcow):

not for exact value....calc gives a decimal approximation

OpenStudy (anonymous):

Yes very true.

OpenStudy (anonymous):

It needs to be in exact terms, i.e. sqrt(3)/2, sqrt(2)/2 etc. This is what I have so far: \[(-1 - \sqrt{3}/3)/1 - \sqrt{3}/3\]

OpenStudy (dumbcow):

 \[\tan(285) = \tan(105) = \tan(60+45)\] \[\tan(a+b) = \frac{\tan a + \tan b}{1-\tan a \tan b}\]

OpenStudy (anonymous):

Ah I was doing tan(315 - 30) which might've been more complicated than 60 - 45

OpenStudy (dumbcow):

possibly, that will still work finding tan(315) on unit circle is just as easy

OpenStudy (anonymous):

I have sqrt{3} + 1 / 1 - sqrt{3}, where do I go from there?

OpenStudy (dumbcow):

multiply top/bottom by conjugate (1+sqrt3)

OpenStudy (dumbcow):

\[= \frac{4+2\sqrt{3}}{-2} = -2 - \sqrt{3}\]